A horizontal pipe widens from a radius of 0.150 m to 0.280 m. If the speed of the water is 1.60 m/s in the larger pipe, what is the speed in the smaller pipe (in m/s)?
Answer #1

Radius of the larger pipe = R1 = 0.28 m

Cross-sectional area of the larger pipe = A1 = R12

Speed of the water in the larger pipe = V1 = 1.6 m/s

Radius of the smaller pipe = R2 = 0.15 m

Cross-sectional area of the smaller pipe = A2 = R22

Speed of the water in the smaller pipe = V2

By Continuity Equation,

V2 = 5.575 m/s

Speed of the pipe in the smaller pipe = 5.575 m/s
Answer #2

Calculation:

  1. Identify the given values:

    • Radius of smaller pipe, r1=0.150m

    • Radius of larger pipe, r2=0.280m

    • Speed in larger pipe, v2=1.60m/s

  2. Calculate the cross-sectional areas:

    • Area of smaller pipe:

      A1=πr12=π(0.150)2=0.0707m2

    • Area of larger pipe:

      A2=πr22=π(0.280)2=0.246m2

  3. Apply the continuity equation (A1v1=A2v2):

    v1=A2v2A1=(0.246m2)(1.60m/s)0.0707m2v1=5.57m/s

The speed in the smaller pipe is 5.57 m/s


source: Self
answered by: Harshwardhan kunal